OpenStudy (swaqqout_kid1):

Starla bought cloth napkins at an outlet store for $4.00 each, and she also used a $5 off coupon. In the equation below, x represents the number of napkins, and y represents the cost before tax. y = $4.00x - $5 If Starla paid $71.00 before tax, how many napkins did she buy? 18 14 19 20 @ttb123456789

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

71=4.00x-5 71+5=4x x=76/4=19.0

OpenStudy (anonymous):

19

OpenStudy (anonymous):

No direct answers. @ttb123456789

OpenStudy (anonymous):

well i showed explanation what?

OpenStudy (anonymous):

I doesn't matter. The person asking the question can just not look at the explanation and take the answer.

OpenStudy (anonymous):

in COC it is wtitten not to give ONLY answers. provide explanatuion

OpenStudy (anonymous):

Look at what this moderator said:

OpenStudy (anonymous):

ok then